OpenStudy (anonymous):

Radicals with odd index roots (3, 5, 7, etc.) can have: Select one: a. Positive answers only b. Negative answers only c. Both positive and negative answers d. Only odd root answers

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@XVasHappeninZen @myininaya @DangerousJesse

OpenStudy (anonymous):

let me see...

OpenStudy (anonymous):

With cube and other odd roots, you can take them of negative numbers. For example the cube root of 8 is 2 since 2^3 = 2*2*2 = 8. What if it's the cube root of -8? Note that (-2)^3 = (-2)*(-2)*(-2) = -8. When you multiply an odd number of negatives, the answer is negative. That's why you can take odd roots of negative numbers. Does this answer your question or do you know this?

OpenStudy (anonymous):

so the answer would be B negative answers only?

OpenStudy (anonymous):

@XVasHappeninZen

OpenStudy (anonymous):

i think yes

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@DangerousJesse @iambatman

OpenStudy (anonymous):

@igreen help

OpenStudy (anonymous):

should i just go with B

OpenStudy (dangerousjesse):

Sorry, I was dealing with something. One sec.

OpenStudy (anonymous):

no just look at the question

OpenStudy (dangerousjesse):

I'd say C, actually. Because \(^3\sqrt{x}\) and \(^3\sqrt{-x}\) can have odd and even answers, I don't see how -2\(^3\) was relevant, but okay.

OpenStudy (dangerousjesse):

@prowrestler

OpenStudy (anonymous):

\[\sqrt[3]{x^{3}}- 3x^{2}+1=x\]

OpenStudy (dangerousjesse):

This one is going to take too long to type out using Mathjax, so here.